# Spherical maximal functions and Hardy spaces for Fourier integral operators

### v2

@@ 4.1. Fourier integral operators
for $f\in\mathcal{S}(\mathbb{R}^{n})$ and $z\in\mathbb{R}^{n+1}$, and $T_{t}f(x):=Tf(x,t)$. Then, cf. [17], $T$ is a Fourier integral operator of order $m-1/4$, and each $T_{t}$ is a Fourier integral operator of order $m$ associated with a local canonical graph.
To prove sharpness of the results in Section 3.1, we will rely on the following lemma, an improvement of [20, Lemma 7.1].
@@
The proof follows that of [20, Lemma 7.1], but ends with an additional step.
@@
Here we used Taylor approximation and the assumptions that $m=0$ and that $a(0,\eta)=1$ for all $\eta\in\mathrm{supp}(\widehat{h})$.
@@
$\delta(h,\nu,z):=\sup\{|\eta\cdot(\partial_{\eta}\Phi(z,\nu)-\partial_{\eta}\Phi(z,\hat{\eta}))|\mid\eta\in\mathrm{supp}(\widehat{h})\}.$
Here we also used that $|e^{ix}-1|=|\cos(x)-1+i\sin(x)|\leq\sqrt{2}x$ for all $x\in\mathbb{R}$.
It thus remains to show that $\delta(h,\nu,z)\lesssim|z|\gamma(h,\nu)$. To this end, for $\eta\in\mathrm{supp}(\widehat{h})$, we apply another Taylor expansion and write $\partial_{\eta}\Phi(z,\nu)-\partial_{\eta}\Phi(z,\hat{\eta})$ as
@@
Moreover, $\partial_{\eta_{i}}\Phi(z,\eta)$ is positively homogeneous of order zero in $\eta$ for every $1\leq i\leq n$, because $\Phi(z,\eta)$ is positively homogeneous of order one in $\eta$. In particular, the radial derivative $(\eta\cdot\partial_{\eta})\partial_{\eta_{i}}\Phi(z,\hat{\eta})$ vanishes for every $1\leq i\leq n$. Hence
@@
for any $t_{0}\in\mathbb{R}$ and $f\in\mathcal{S}(\mathbb{R}^{n})$, and $d(p)+\frac{1}{p}=s(p)=\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})$ for $p\geq\frac{2(n+1)}{n-1}$. Hence the following result shows that, under mild conditions on the symbol of the operator, the maximal function bounds in Theorems 3.1 and 3.4, and in Corollary 3.2 and Remark 3.3, are essentially sharp for such $p$.
@@
The proof is analogous to that of [20, Theorem 7.3]. However, given that the setting under consideration presently is slightly different, and because we will use some of the ideas later on, we provide most of the details.
@@
After these reduction steps, we get to the heart of the proof. For each $k\in\mathbb{Z}_{+}$, let $\Theta_{k}\subseteq S^{n-1}$ be a maximal collection of unit vectors such that $|\nu-\nu^{\prime}|\geq 2^{-k/2}$ for all $\nu,\nu^{\prime}\in\Theta_{k}$ with $\nu\neq\nu^{\prime}$. Let $V^{\prime\prime}\subseteq V^{\prime}$ be a non-empty cone which is closed in $V^{\prime}$, and fix a $k\in\mathbb{Z}_{+}$ such that $V_{k}:=\Theta_{k}\cap V^{\prime\prime}$ has approximately $2^{k(n-1)/2}$ elements. Note that this is true for all large $k$, since $V^{\prime\prime}$ has a nonzero aperture. Next, let $(\chi_{\nu})_{\nu\in\Theta_{k}}\subseteq C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be an associated partition of unity. In particular, each $\chi_{\nu}$ is positively homogeneous of degree 00 and satisfies $0\leq\chi_{\nu}\leq 1$,
@@
with the obvious modification for $p=\infty$, as follows from [27, Proposition 4.1].
Next, note that $\Phi$ and all of its derivatives are bounded on compact subsets of $\mathbb{R}^{n+1}\times(\mathbb{R}^{n}\setminus\{0\})$. Hence $z\mapsto\partial_{\eta}\Phi(z,\nu)$ is a locally Lipschitz flow for each $\nu\in V_{k}$, with Lipschitz constants independent of $\nu$. Moreover, by assumption, $\partial_{\eta}\Phi(0,\nu)=0$ and $f_{\nu}(0)=\|\widehat{\psi}\|_{L^{1}(\mathbb{R}^{n})}>0$ for each $\nu$. It thus follows that $\operatorname{Re}(f_{\nu}(\partial_{\eta}\Phi(x,0,\nu)))\gtrsim 1$ whenever $|x|\lesssim 2^{-k}$ for some implicit constant independent of $k$ and $\nu$. On the other hand, $|\hat{\eta}-\nu|^{2}|\eta|\lesssim 1$ for all $\eta\in\mathrm{supp}(\widehat{f_{\nu}}\,)$. Hence Lemma 4.1 yields
@@
Note that Theorem 3.1 is in fact a statement about a single Fourier integral operator acting on function spaces over $\mathbb{R}^{n}$, and the implicit time parameter plays no role in the proof. Combined with the Sobolev embeddings in (2.2), the theorem shows that the bounds in Proposition 2.2 (1) and Proposition 2.3 (1) are sharp under very mild assumptions, for $p\in[2,\infty)$. It was already shown in [27, 20] that Proposition (2.2) (2) and Proposition 2.3 (2) are sharp under the same assumptions.
@@
The conclusion of Theorem 4.2 essentially follows from [30] (see also [33, pp. 185-186]) if the projection of the canonical relation
@@
is bounded. This assumption on the rank of the projection is satisfied in the setting of Theorem 1.2 and, combined with some of the arguments later on in this section, also in the results in Sections 3.2, 3.3 and 3.4. On the other hand, (4.5) definitely does not imply that $s\geq s(p)$ for general $\Phi$ as in Theorem 4.2, since the latter result also applies when, for example, $T_{t_{0}}$ is the identity operator.
@@
We will now show that nontrivial dependence of the phase on $t$ is in fact the only additional assumption required to prove the desired sharpness bound. The following result implies in particular that, under mild conditions on the symbol of the operator, the maximal function bounds in Theorems 3.1 and 3.4, as well as in Corollary 3.2 and Remark 3.3, are essentially sharp for $1\leq p\leq 2$.
@@
by [27, Proposition 2.2]. Here the implicit constants depend on $\theta$, but this does not matter since we will only apply (4.7) once we have already fixed $\theta$. Finally, there exist $c_{0},\delta>0$ such that
@@
$-e_{1}\cdot\gamma_{0}^{\prime}(0)=e_{1}\cdot\partial_{t}\partial_{\eta}\Phi(0,0,e_{1})=\partial_{t}(e_{1}\cdot\partial_{\eta}\Phi(0,0,e_{1}))=\partial_{t}\Phi(0,0,e_{1})\neq 0.$
@@
Then, since the $(n-1)$-dimensional measure of $E_{0}$ is approximately $2^{-k\frac{n-1}{2}}$, the $n$-dimensional measure of $E$ is approximately $2^{-k\frac{n-1}{2}}$ as well. Again, here the implicit constants depend on $\theta$, but this will not matter once $\theta$ has been fixed.
@@
Recall that, under the assumptions of Theorems 3.1 and 3.4, the Sobolev embeddings in (2.2) yield an inequality of the form
@@ 4.2. Sharpness for hypersurfaces
We first prove sharpness in the case of Theorem 3.6. Recall the definitions of $\mathcal{A}_{t}$ and $\mathcal{M}_{\Sigma,\psi}$ from (3.11) and (3.12), for $\Sigma\subseteq\mathbb{R}^{n}$ a hypersurface, $\psi\in C^{\infty}_{c}(\Sigma)$ and $t>0$. The following result shows that the maximal function estimates in Theorem 3.6 are essentially sharp for $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$ in all nontrivial cases, and in particular in the case of Theorem 1.1.
@@
Let $p\in[1,\infty]$ and $s\in\mathbb{R}$. Let $\Sigma\subseteq\mathbb{R}^{n}$ be a hypersurface with non-vanishing Gaussian curvature, and let $\psi\in C^{\infty}_{c}(\Sigma)$ not be identically zero. If there exist $t_{0},C_{t_{0}}>0$ such that
@@
We would like to apply Theorems 4.2 and 4.5. This is complicated by the fact that, although one may express $\mathcal{A}_{t}$ in terms of a sum of operators as in (4.1), the contributions from the summands may, a priori, cancel each other out. To deal with this issue we will microlocalize in regions of phase space where there is only a single summand. To this end we will reason in a manner which can also be used for other results in this section. Throughout, we freely make use of the theory of Fourier integral operators as can be found in [17, Chapter XXV] (see also [33, Chapter 6]).
We first do some preliminary work. Note that we can represent $\mathcal{A}_{1}$ as in (3.19):
@@
for all $t>0$, $f\in\mathcal{S}(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$, where $\Psi_{t}(x,y)=\Psi_{1}(x/t,y/t)$ and $\rho_{0}(x,t,y)=t^{-n}\rho_{0}(x/t,1,y/t)$. Moreover, the assumption that $\Sigma$ has non-vanishing Gaussian curvature is equivalent to (3.15) (see e.g. [36, Section XI.3.1]), and $\rho_{0}(\cdot,t,\cdot)$ is not identically zero. It follows that each $\mathcal{A}_{t}$ is a Fourier integral operator of order $-(n-1)/2$ with canonical relation
@@
$(y,\theta)\mapsto\big{(}\partial_{x}(\theta\Psi_{t}(x,y)),\partial_{\theta}(\theta\Psi_{t}(x,y))\big{)}$
does not vanish where $\Psi_{t}(x,y)=0$ and $\theta=1$. Hence one can solve $\Psi_{t}(x,y)=0$ and $\theta\partial_{x}\Psi_{t}(x,y)=\xi$ with respect to $(y,\theta)$ and use $(x,\xi)$ as local coordinates on $\mathcal{C}_{t}$. This in turn implies that the projections from $\mathcal{C}_{t}$ onto the $x$ coordinate and onto the last two coordinates are submersions. Finally, since $\rho_{0}(\cdot,t,\cdot)$ is not identically zero, the principal symbol of $\mathcal{A}_{t}$ does not vanish at some point $c_{t}=(x_{t},\xi_{t},y_{t},\eta_{t})\in\mathcal{C}_{t}$.
Now suppose that (4.11) holds. Let $S_{1}$ and $S_{2}$ be pseudodifferential operators of order zero, with principal symbols that are nonzero at $(x_{t_{0}},\xi_{t_{0}})$ and $(y_{t_{0}},\eta_{t_{0}})$, respectively, and with microsupport contained in small conic neighborhoods of these points. Since both $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ and $L^{p}(\mathbb{R}^{n})$ are invariant under pseudodifferential operators of order zero, one has
$\|S_{1}\mathcal{A}_{t_{0}}S_{2}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Moreover, the canonical relation of $S_{1}\mathcal{A}_{t_{0}}S_{2}$ is contained in a small conic neighborhood of $c_{t_{0}}$, and its principal symbol does not vanish at $c_{t_{0}}$. In fact, after multiplying by a smooth cutoff, we may additionally suppose that $S_{1}\mathcal{A}_{t_{0}}S_{2}f=0$ if $f$ is supported outside a small neighborhood of $y_{t_{0}}$.
It then follows (see [20, Lemma A.2 and Remark A.4]) that one can write $S_{1}\mathcal{A}_{t_{0}}S_{2}=T_{t_{0}}S_{t_{0}}+R_{t_{0}}$, where $R_{t_{0}}$ is a smoothing operator, $S_{t_{0}}$ is a change of coordinates and $T_{t_{0}}$ is an operator as in (4.1), with $m=-(n-1)/2$, albeit with the $(n+1)$-dimensional $z$ variable replaced by the $n$-dimensional $x$ variable. Moreover, $T_{t_{0}}$ satisfies the $n$-dimensional version of the conditions in Theorem 4.2. Here we have used the projection conditions on $\mathcal{C}_{t_{0}}$.
Since $\|S_{t_{0}}f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}\eqsim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for $f\in\mathcal{S}(\mathbb{R}^{n})$ supported in a suitably small neighborhood of $y_{t_{0}}$, and because $R_{t_{0}}:\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})\to L^{p}(\mathbb{R}^{n})$, we have
@@
Next, suppose that (4.12) holds. We will reduce to Theorem 4.5 in a similar manner as above, with the minor caveat that the mixed norm implicit on the left-hand side of (4.12) is not invariant under applications of general pseudodifferential operators, and we need to be more careful when microlocalizing. More importantly, we have to work with a phase function satisfying the assumptions of Theorem 4.5.
@@
Set $\mathcal{A}f(x,t):=\chi(t)\mathcal{A}_{t}f(x)$ for $f\in\mathcal{S}(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$, where $\chi\in C^{\infty}_{c}(\mathbb{R})$ satisfies $\mathrm{supp}(\chi)\subseteq[1,2]$ and does not vanish at some point $t_{0}\in(1,2)$. Then $\mathcal{A}$ is a Fourier integral operator of order $-\frac{n-1}{2}-\frac{1}{4}$ with canonical relation
(4.14) $\mathcal{C}:=\{(x,t,\theta\partial_{x}\Psi_{t}(x,y),\theta\partial_{t}\Psi_{t}(x,y),y,\theta\partial_{x}\Psi_{t}(x,y))\mid y\in x-\Sigma,\theta\in\mathbb{R}\setminus\{0\}\}.$
@@
As before, the projections onto the $(x,t)$ coordinate and onto the last two coordinates are submersions. This in turn implies that $\mathcal{C}$ can be locally parametrized using the $(x,t)$ coordinate and the last coordinate. Moreover, the function $\chi(t_{0})\rho_{0}(\cdot,t_{0},\cdot)$ is not identically zero. Hence the principal symbol of $\mathcal{A}$ does not vanish at some point $c_{0}=(x_{0},t_{0},\xi_{0},\tau_{0},y_{0},\eta_{0})\in\mathcal{C}$, and we may suppose that $x_{0}\neq y_{0}$.
@@
Let $S_{2}$ be a pseudodifferential operator of order zero with a principal symbol which is nonzero at $(y_{0},\eta_{0})$, and with microsupport contained in a small conic neighborhood of this point. Let $\chi_{1}\in C^{\infty}_{c}(\mathbb{R}^{n})$ be nonzero at $(x_{0},t_{0})$ and supported in a small neighborhood of this point. Then
$\big{\|}\sup_{t\in\mathbb{R}}|\chi_{1}(\cdot,t)\mathcal{A}S_{2}f(\cdot,t)|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\big{\|}\sup_{1\leq t\leq 2}|\chi(t)\mathcal{A}_{t}S_{2}f|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
@@
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Here we used (4.12) and that $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under $S_{2}$. Moreover, since $\mathcal{C}$ can be parametrized using the $(x,t)$ coordinate and its last coordinate, it suffices to localize near $(x_{0},t_{0})$ and microlocalize near $(y_{0},\eta_{0})$, as we have done, to guarantee that the canonical relation of $\chi_{1}\mathcal{A}S_{2}$ is contained in a small conical neighborhood of $c_{0}$. Finally, after multiplying by a smooth cutoff, we may suppose that $\chi_{1}\mathcal{A}S_{2}f=0$ if $f$ is supported outside a small neighborhood of $y_{0}$.
@@
It then follows (see [20, Lemma A.2 and Remark A.4]) that $\chi_{1}\mathcal{A}S_{2}=TS+R$, where $R$ is a smoothing operator, $S$ is a change of coordinates and $T$ is an operator as in (4.1), with $m=-(n-1)/2$, satisfying the conditions of Theorem 4.2. One again has $\|Sf\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}\eqsim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for $f$ supported in a suitably small neighborhood of $y_{0}$. Moreover, as in the proof of Theorem 3.1, one can show that $\|\sup_{t\in\mathbb{R}}|Rf(\cdot,t)|\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Hence

### v3

@@ 4.1. Fourier integral operators
for $f\in\mathcal{S}(\mathbb{R}^{n})$ and $z\in\mathbb{R}^{n+1}$, and $T_{t}f(x):=Tf(x,t)$. Then, cf. [19], $T$ is a Fourier integral operator of order m1/4m-1/4, and each $T_{t}$ is a Fourier integral operator of order $m$ associated with a local canonical graph.
To prove sharpness of the results in Section 3.1, we will rely on the following lemma, an improvement of [22, Lemma 7.1].
@@
The proof follows that of [22, Lemma 7.1], but ends with an additional step.
@@
[…]
@@
[…]
Here we also used that $|e^{ix}-1|=|\cos(x)-1+i\sin(x)|\leq\sqrt{2}|x|$ for all $x\in\mathbb{R}$.
[…]
@@
[…]
@@
for any $t_{0}\in\mathbb{R}$ and $f\in\mathcal{S}(\mathbb{R}^{n})$, and $d(p)+\frac{1}{p}=s(p)=\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})$ for $p\geq\frac{2(n+1)}{n-1}$. Hence the following result shows that, under mild conditions on the symbol of the operator, the maximal function bounds in Theorems 3.2 and 3.5, and in Corollary 3.3 and Remark 3.4, are essentially sharp for such $p$.
@@
The proof is analogous to that of [22, Theorem 7.3]. However, given that the setting under consideration presently is slightly different, and because we will use some of the ideas later on, we provide most of the details.
@@
After these reduction steps, we get to the heart of the proof. For each $k\in\mathbb{Z}_{+}$, let $\Theta_{k}\subseteq S^{n-1}$ be a maximal collection of unit vectors such that $|\nu-\nu^{\prime}|\geq 2^{-k/2}$ for all $\nu,\nu^{\prime}\in\Theta_{k}$ with $\nu\neq\nu^{\prime}$. Let $V^{\prime\prime}\subseteq V^{\prime}$ be a non-empty cone which is closed in $V^{\prime}$, and fix a $k\in\mathbb{Z}_{+}$ such that $V_{k}:=\Theta_{k}\cap V^{\prime\prime}$ has approximately $2^{k(n-1)/2}$ elements. Note that this is true for all large $k$, since $V^{\prime\prime}$ has a nonzero aperture. Next, let $(\chi_{\nu})_{\nu\in\Theta_{k}}\subseteq C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be an associated partition of unity. In particular, each $\chi_{\nu}$ is positively homogeneous of degree 0 and satisfies $0\leq\chi_{\nu}\leq 1$,
@@
with the obvious modification for $p=\infty$, as follows from [31, Proposition 4.1].
[…]
@@
Note that Theorem 3.2 is in fact a statement about a single Fourier integral operator acting on function spaces over $\mathbb{R}^{n}$, and the implicit time parameter plays no role in the proof. Combined with the Sobolev embeddings in (2.2), the theorem shows that the bounds in Proposition 2.2 (1) and Proposition 2.3 (1) are sharp under very mild assumptions, for $p\in[2,\infty)$. It was already shown in [31, 22] that Proposition 2.2 (2) and Proposition 2.3 (2) are sharp under the same assumptions.
@@
The conclusion of Theorem 4.2 essentially follows from [34] (see also [37, pp. 185-186]) if the projection of the canonical relation
@@
is bounded. This assumption on the rank of the projection is satisfied in the setting of Theorem 1.2 and (1.9), and, combined with some of the arguments later on in this section, also in (1.8) and in the results in Sections 3.2, 3.3 and 3.4. On the other hand, (4.5) definitely does not imply that $s\geq s(p)$ for general $\Phi$ as in Theorem 4.2, since the latter result also applies when, for example, $T_{t_{0}}$ is the identity operator.
@@
We will now show that nontrivial dependence of the phase on $t$ is in fact the only additional assumption required to prove the desired sharpness bound. The following result implies in particular that, under mild conditions on the symbol of the operator, the maximal function bounds in Theorems 3.2 and 3.5, as well as in Corollary 3.3, are essentially sharp for $1\leq p\leq 2$.
@@
by [31, Proposition 2.2]. Here the implicit constants depend on $\theta$, but this does not matter since we will only apply (4.7) once we have already fixed $\theta$. Finally, there exist $c_{0},\delta>0$ such that
@@
$e_{1}\cdot\gamma_{0}^{\prime}(0)=e_{1}\cdot\partial_{t}\partial_{\eta}\Phi(0,0,e_{1})=e_{1}\cdot\partial_{\eta}(\partial_{t}\Phi)(0,0,e_{1})=\partial_{t}\Phi(0,0,e_{1})\neq 0.$
@@
Then, since the (n1)(n-1)-dimensional measure of $E_{0}$ is approximately $2^{-k\frac{n-1}{2}}$, the $n$-dimensional measure of $E$ is approximately $2^{-k\frac{n-1}{2}}$ as well. Again, here the implicit constants depend on $\theta$, but this will not matter once $\theta$ has been fixed.
@@
Recall that, under the assumptions of Theorems 3.2 and 3.5, the Sobolev embeddings in (2.2) yield an inequality of the form
@@ 4.2. Sharpness for hypersurfaces
We first prove sharpness in the case of Theorem 3.7. Recall the definitions of $\mathcal{A}_{t}$ and $\mathcal{M}_{\Sigma,\psi}$ from (3.11) and (3.12), for $\Sigma\subseteq\mathbb{R}^{n}$ a hypersurface, $\psi\in C^{\infty}_{c}(\mathbb{R}^{n})$ and $t>0$. The following result shows that the maximal function estimates in Theorem 3.7 are essentially sharp for $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ in all nontrivial cases, and in particular in the case of Theorem 1.1 (see also Remark 3.4).
@@
Let $p\in[1,\infty]$ and $s\in\mathbb{R}$. Let $\Sigma\subseteq\mathbb{R}^{n}$ be a hypersurface with non-vanishing Gaussian curvature, and let $\psi\in C^{\infty}_{c}(\mathbb{R}^{n})$ not be identically zero. If there exist $t_{0},C_{t_{0}}>0$ such that
@@
We would like to apply Theorems 4.2 and 4.5. This is complicated by the fact that, although one may express $\mathcal{A}_{t}$ in terms of a sum of operators as in (4.1), the contributions from the summands may, a priori, cancel each other out. To deal with this issue we will microlocalize in regions of phase space where there is only a single summand. We will reason in a manner which can also be used for other results in this section. Throughout, we freely make use of the theory of Fourier integral operators as can be found in [19, Chapter XXV] (see also [37, Chapter 6]).
Preliminary work
Note that we can represent $\mathcal{A}_{1}$ as in (3.22):
@@
for all $t>0$, $f\in\mathcal{S}(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$, where $\Psi_{t}(x,y)=\Psi_{1}(x/t,y/t)$ and $\rho_{0}(x,t,y)=t^{-n}\rho_{0}(x/t,1,y/t)$. Moreover, the assumption that $\Sigma$ has non-vanishing Gaussian curvature is equivalent to (3.18) (see e.g. [40, Section XI.3.1]), and $\rho_{0}(\cdot,t,\cdot)$ is not identically zero. It follows that each $\mathcal{A}_{t}$ is a Fourier integral operator of order (n1)/2-(n-1)/2 with canonical relation
@@
$(y,\theta)\mapsto\big(\partial_{x}(\theta\Psi_{t}(x,y)),\theta\Psi_{t}(x,y)\big)$
does not vanish where $\Psi_{t}(x,y)=0$ and $\theta\neq 0$. Hence one can solve $\Psi_{t}(x,y)=0$ and $\theta\partial_{x}\Psi_{t}(x,y)=\xi$ with respect to $(y,\theta)$ and use $(x,\xi)$ as local coordinates on $\mathcal{C}_{t}$. This in turn implies that the projection from $\mathcal{C}_{t}$ onto the $(x,\xi)$ coordinates is a submersion. Similarly, one can use $(y,\eta)$ as local coordinates on $\mathcal{C}_{t}$, and the projection from $\mathcal{C}_{t}$ onto the $(y,\eta)$ coordinates is also a submersion.
Finally, since $\rho_{0}(\cdot,t,\cdot)$ is not identically zero, the principal symbol of $\mathcal{A}_{t}$ does not vanish at some point $c_{t}=(x_{t},\xi_{t},y_{t},\eta_{t})\in\mathcal{C}_{t}$.
Proof of the first statement
Suppose that (4.11) holds. We will microlocalize by multiplying $\mathcal{A}_{t_{0}}$ from the right by a pseudodifferential operator. More precisely, let $S$ be a pseudodifferential operator of order zero on $\mathbb{R}^{n}$, with a principal symbol which is nonzero at $(y_{t_{0}},\eta_{t_{0}})$, and with microsupport contained in a small conic neighborhood of this point. Since $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under pseudodifferential operators of order zero, one then has
$\|\mathcal{A}_{t_{0}}Sf\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. As noted above, the canonical relation of $\mathcal{A}_{t_{0}}$, which in turn equals that of $\mathcal{A}_{t_{0}}S$, can be locally parametrized near $c_{t_{0}}=(x_{t_{0}},\xi_{t_{0}},y_{t_{0}},\eta_{t_{0}})$ using the $(y,\eta)$ coordinates. Given that $S$ has microsupport near $(y_{t_{0}},\eta_{t_{0}})$, the canonical relation of $\mathcal{A}_{t_{0}}S$ is thus contained in a small conic neighborhood of $c_{t_{0}}$, and its principal symbol does not vanish at $c_{t_{0}}$. In fact, after multiplying on the right by a smooth cutoff, we may additionally suppose that $\mathcal{A}_{t_{0}}Sf=0$ if $f$ is supported outside a small neighborhood of $y_{t_{0}}$.
It then follows (see [22, Lemma A.2 and Remark A.4]) that one can write $\mathcal{A}_{t_{0}}S=T_{t_{0}}S_{t_{0}}^{\prime}+R_{t_{0}}$, where $R_{t_{0}}$ is a smoothing operator, $S_{t_{0}}^{\prime}$ is a change of coordinates and $T_{t_{0}}$ is an operator as in (4.1), with m=(n1)/2m=-(n-1)/2, albeit with the $(n+1)$-dimensional $z$ variable replaced by the $n$-dimensional $x$ variable. Moreover, $T_{t_{0}}$ satisfies the $n$-dimensional version of the conditions in Theorem 4.2. Here we have used the projection conditions on $\mathcal{C}_{t_{0}}$.
Since $\|{S_{t_{0}}^{\prime}}f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}\eqsim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for $f\in\mathcal{S}(\mathbb{R}^{n})$ supported in a suitably small neighborhood of $y_{t_{0}}$, and because $R_{t_{0}}:\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})\to L^{p}(\mathbb{R}^{n})$, we have
@@
More preliminary work
We will now prepare for the proof of the second statement.
@@
Set $\mathcal{A}f(x,t):=\chi(t)\mathcal{A}_{t}f(x)$ for $f\in\mathcal{S}(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$, where $\chi\in C^{\infty}_{c}(\mathbb{R})$ satisfies $\mathrm{supp}(\chi)\subseteq[1,2]$ and does not vanish at some point $t_{0}\in(1,2)$. Then $\mathcal{A}$ is a Fourier integral operator of order $\frac{n-1}{2}-\frac{1}{4}$ with canonical relation
(4.14) $\mathcal{C}:=\{(x,t,\theta\partial_{x}\Psi_{t}(x,y),\theta\partial_{t}\Psi_{t}(x,y),y,\theta\partial_{y}\Psi_{t}(x,y))\mid y\in x-\Sigma,\theta\in\mathbb{R}\setminus\{0\}\}.$
@@
The projections onto the $(x,t)$ coordinate and onto the last two coordinates are submersions. This implies that, if the $(x,t)$ coordinate and the last two coordinates are fixed on $\mathcal{C}$, then so are the remaining coordinates. Moreover, the function $\chi(t_{0})\rho_{0}(\cdot,t_{0},\cdot)$ is not identically zero. Hence the principal symbol of $\mathcal{A}$ does not vanish at some point $c_{0}=(x_{0},t_{0},\xi_{0},\tau_{0},y_{0},\eta_{0})\in\mathcal{C}$. In fact, the principal symbol is nonzero in an open neighborhood of $c_{0}$ in $\mathcal{C}$, and the $y$-coordinate in this neighborhood may vary over an open neighborhood within the hypersurface $x_{0}-\Sigma$, by (4.14). In particular, we may thus assume that $x_{0}\neq y_{0}$.
Proof of the second statement
Suppose that (4.12) holds. As in the proof of the first statement, we will microlocalize $\mathcal{A}$. However, unlike in the setting of the first statement, $\mathcal{C}$ cannot be locally parametrized using only its last two coordinates, so we also need to multiply $\mathcal{A}$ on the left by a pseudodifferential operator. On the other hand, the mixed norm implicit on the left-hand side of (4.12) is not invariant under general pseudodifferential operators, so we have to be slightly careful here.
@@
Again let $S$ be a pseudodifferential operator of order zero with a principal symbol which is nonzero at $(y_{0},\eta_{0})$, and with microsupport contained in a small conic neighborhood of this point. Let $\chi_{1}\in C^{\infty}_{c}(\mathbb{R}^{n+1})$ be nonzero at $(x_{0},t_{0})$ and supported in a small neighborhood of this point. Then
$\big\|\sup_{t\in\mathbb{R}}|\chi_{1}(\cdot,t)\mathcal{A}Sf(\cdot,t)|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\big\|\sup_{1\leq t\leq 2}|\chi(t)\mathcal{A}_{t}Sf|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
@@
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Here we used (4.12), that $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under $S$, and that $\chi$ and $\chi_{1}$ are multiplication operators. Moreover, since $\mathcal{C}$ is determined by its $(x,t)$ coordinate and its last two coordinates, it suffices to localize near $(x_{0},t_{0})$ and microlocalize near $(y_{0},\eta_{0})$, as we have done, to guarantee that the canonical relation of $\chi_{1}\mathcal{A}S$ is contained in a small conic neighborhood of $c_{0}$. Finally, after multiplying by another smooth cutoff, we may suppose that $\chi_{1}\mathcal{A}Sf=0$ if $f$ is supported outside a small neighborhood of $y_{0}$.
@@
It then follows (see [22, Lemma A.2 and Remark A.4]) that $\chi_{1}\mathcal{A}S=TS^{\prime}+R$, where $R$ is a smoothing operator, $S^{\prime}$ is a change of coordinates and $T$ is an operator as in (4.1), with m=(n1)/2m=-(n-1)/2, satisfying the conditions of Theorem 4.2. One again has $\|{S^{\prime}}f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}\eqsim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for $f$ supported in a suitably small neighborhood of $y_{0}$. Moreover, as in the proof of Theorem 3.2, one can show that $\|\sup_{t\in\mathbb{R}}|Rf(\cdot,t)|\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Hence